Servers status
Check our services status at a glance
- ID:
- 20
- Title:
- http7 hardware replacement
- Status:
- completed
- Started date:
- 01/22/2012 midnight
- End date:
- 01/22/2012 12:20 a.m.
- Involved servers:
- http7
Upgrades
01/21/2012
15:55
15:55
http7 disk cables and /dev/sda will be replaced tonight. The server will be unavailable during the operation.
01/22/2012
0:03
0:03
The operation has started.
01/22/2012
0:09
0:09
The operation completed successfully.
01/26/2012
11:36
11:36
The underlying issue (complete disks lockup despite RAID1) is still present. /dev/sdb will be replaced tonight from 00:00 (CET). Expected downtime: less than 10 minutes.
01/27/2012
0:07
0:07
The operation has started.
01/27/2012
0:11
0:11
The operation completed successfully. Total downtime: 5 minutes.